OpenStudy (anonymous):

what is the domain of sqr, root 3x + 15

OpenStudy (anonymous):

\[\sqrt{3x+15}\] You know that has to be greater than or equal to 0 because square roots are always positive. You can solve an inequality the same way that you would solve any other equation, in this case, by isolating the X :)

OpenStudy (anonymous):

\[\sqrt{3x+15 what is the domain}\]

OpenStudy (anonymous):

so positve integer to infinity ?

OpenStudy (anonymous):

I got \[x \ge -5\] because the expression under the radical sign needs to be positive. That's the only restriction :)